OpenStudy (anonymous):

can someone check my answer? What is the least common denominator of the fraction 2 over x plus the fraction x over x plus 1 plus the fraction x plus 1 over 3 x ? 3x + 1 3(x2 + 1) 3(x + 3) 3x(x + 1) < --- my choice

OpenStudy (anonymous):

\[\frac{ 2 }{ x } + \frac{ x }{ x + 1 } + \frac{ x + 1 }{ 3x }\]

OpenStudy (anonymous):

lol x

OpenStudy (anonymous):

least common denominator not numerator

OpenStudy (anonymous):

so something that can multiply into all of the bottom =)

OpenStudy (anonymous):

i think its 3x(x + 1)

OpenStudy (anonymous):

heres the definition "another term for lowest common denominator".

OpenStudy (anonymous):

yeah i know what denominator means its just that when i see a variable it confuses me

OpenStudy (anonymous):

A =)

OpenStudy (anonymous):

3x x x gives you 3x^2 i believe =)

OpenStudy (anonymous):

i thought 3x x x = 3x^3 because you have 3 x's

OpenStudy (anonymous):

so you don't think its D you think its A?

OpenStudy (anonymous):

ur right =)

OpenStudy (anonymous):

lol i was confusing myself ... idk how though ...

OpenStudy (anonymous):

so it's D ?

OpenStudy (anonymous):

Yeah ...

OpenStudy (anonymous):

k thnx

OpenStudy (anonymous):

yeah sorry i was thinking of some other stupid formula ...

OpenStudy (anonymous):

im going to ask @dan815 and @nincompoop @Preetha to check my work

OpenStudy (anonymous):

tbh thats why i hate math cuz when you learn a formula they teach you a new one and then you get recked and have problems and then people shoot up schools

OpenStudy (anonymous):

here =)

OpenStudy (anonymous):

is fine

OpenStudy (anonymous):

look at it there =)

OpenStudy (anonymous):

@uri d okay give me a sec